A car dealership interviewed every driver who came onto their lot on Tuesday. This produced the following results: there were 31 people who owned trucks and 62 people who owned cars. If 22 people owned both a truck and a car and 9 people owned neither, how many people were interviewed?

Answer by nyc_function(2741)

If 31 people owned a truck and 22 owned both a truck and car, then 31-22 = 9 people owned just a truck. Likewise, 62-22=40 people owned just a car. So the total number of people interviewed was:
(people who own just trucks) + (people who own just cars) + (people who own both) + (people who own neither) =
9 + 40 + 22 + 9 = 80 people